# Embedding Inversion via Conditional Masked Diffusion Language Models

### v2

@@ Abstract
We frame embedding inversion as conditional masked diffusion, recovering all tokens in parallel through iterative denoising rather than sequential autoregressive generation. A masked diffusion language model is conditioned on the target embedding via adaptive layer normalization, requiring only 8 forward passes through a 78M parameter model with no access to the target encoder. On 32-token sequences across three embedding models, the method achieves up to 81.3% token accuracy. Source code and live demo are available at https://github.com/jina-ai/embedding-inversion-demo.
@@ 1 Introduction
We propose an alternative formulation: embedding inversion as conditional masked diffusion. Starting from a fully masked sequence, a denoising model iteratively reveals tokens at all positions in parallel, conditioned on the target embedding vector via adaptive layer normalization. Correction is built into the diffusion process itself: each step refines all positions simultaneously using global context, without ever reembedding the current hypothesis. The embedding vector enters only through AdaLN modulation, making the approach encoder agnostic: the same architecture applies to any embedding model without alignment training. We validate on three encoders with different architectures and dimensionalities, achieving up to 81.3% token recovery on 32-token sequences with no access to the target encoder at inference time.
@@ 3.3 Model Architecture
where $\mathbf{W}_{1}\in\mathbb{R}^{D_{h}\times d}$, $\mathbf{W}_{2}\in\mathbb{R}^{D_{h}\times D_{h}}$, and $\mathbf{b}_{1},\mathbf{b}_{2}\in\mathbb{R}^{D_{h}}$ are learned parameters. We use an 8-layer transformer with 12 attention heads and FFN dimension $4D_{h}$. In our experiments, $D_{h}=768$, yielding FFN dimension 3072. Input and output embeddings are weight-tied to reduce parameters given the large vocabulary size $|\mathcal{V}|=50257$.
@@
where $\text{MLP}_{t}^{(\ell)}$ and $\text{MLP}_{c}^{(\ell)}$ are single-layer MLPs that output vectors of dimension $D_{h}$. The layer normalization at layer $\ell$ is then modulated:
@@ 3.4 Decoding Strategies
We consider four strategies for generating tokens from the trained model.
@@ 4 Experimental Results
We train on 2M samples from C4 (Raffel et al., 2020), filtered to 32 tokens. We use the GPT-2 tokenizer with vocabulary size 50,257. Training uses batch size 400 for 200K steps with AdamW optimizer at learning rate $10^{-4}$ and EMA decay 0.9999. We employ a log-linear noise schedule with $\lambda=5.0$ following Sahoo et al. (2024). Timesteps are sampled uniformly from $[0,1]$. Embeddings are computed using the target encoder and cached. We evaluate on three embedding models with different architectures and dimensionalities: jina-embeddings-v3 (Sturua et al., 2025) with 570M parameters and 1024-dimensional embeddings, Qwen3-Embedding-0.6B with 600M parameters and 1024-dimensional embeddings, and EmbeddingGemma-300m with 300M parameters and 768-dimensional embeddings. We train separate models for each encoder using multilingual data from mC4 to assess generalization across embedding spaces.
The complete model has approximately 270M parameters due to the large vocabulary embeddings, but only 78M trainable parameters consisting of the 8 transformer layers, embedding projection MLP, and AdaLN conditioning MLPs.
Table 1 shows results across all three embedding encoders using sequential greedy decoding, which provides the highest token accuracy. Qwen3-Embedding achieves the best performance at 81.3% token accuracy, followed by EmbeddingGemma at 78.8% and jina-v3 at 76.0%. All models are trained on multilingual data from mC4.
Table 1: Performance across embedding encoders using sequential greedy decoding. All trained on 2M multilingual samples from mC4. Best checkpoint selected by validation loss.
@@
Table 2 compares four decoding strategies across all three encoders on 10 languages. Cosine similarity is averaged over the same sentence translated into English, Chinese, German, Japanese, French, Spanish, Korean, Russian, Arabic, and Portuguese. Sequential greedy consistently achieves the highest similarity across encoders.
@@
Euler with remasking at 0.05 improves over vanilla Euler by 2.6 percentage points in token accuracy. Two-stage decoding achieves highest exact match at 13.1%. Baselines confirm that embedding conditioning is essential: random tokens achieve 0.02% accuracy, while unconditional LM achieves 2.1% despite high fluency with BLEU score 89.3.
@@
Figure 2: Training dynamics across three embedding encoders on 2M multilingual samples. Qwen3-Embedding reaches 81.3% token accuracy at 72.5K steps with validation loss 1.32. All models show diminishing returns beyond 50K steps, suggesting architectural improvements rather than extended training as the path to further gains.
5 Conclusion
We presented embedding inversion via conditional masked diffusion, achieving 81.3% token accuracy across three embedding models with a 78M parameter decoder that requires no access to the target encoder. The progression from Vec2Text, which requires a compatible encoder architecture, to ALGEN, which needs alignment training, to Zero2Text, which queries the target API, to our work, which needs none of these, demonstrates that inversion attacks are becoming more accessible, not less. Production systems that cache or transmit embedding vectors under the assumption of irreversibility should treat them as sensitive data requiring protection equivalent to the original text.
Current limitations include the restriction to 32-token sequences, whereas real documents are substantially longer. Scaling to longer sequences via hierarchical diffusion or sliding window approaches is a natural next step. Incorporating language model priors through classifier-free guidance could further improve reconstruction quality, particularly for low-frequency tokens. On the defense side, encrypted vector search and inversion-resistant embeddings through adversarial training or differential privacy on the embedding space remain important open problems for security-critical deployments.

### v3

@@ Abstract
We frame embedding inversion as conditional masked diffusion, recovering all tokens in parallel through iterative denoising rather than sequential autoregressive generation. A masked diffusion language model is conditioned on the target embedding via adaptive layer normalization, requiring only 8 forward passes with no access to the target encoder at inference time. On 32-token sequences across three embedding models, the method achieves token recovery through parallel denoising without requiring encoder access, iterative correction, or architecture-specific alignment. Source code and live demo are available at https://github.com/jina-ai/embedding-inversion-demo.
@@ 1 Introduction
We propose an alternative formulation: embedding inversion as conditional masked diffusion. Starting from a fully masked sequence, a denoising model iteratively reveals tokens at all positions in parallel, conditioned on the target embedding vector via adaptive layer normalization. Correction is built into the diffusion process itself: each step refines all positions simultaneously using global context, without ever reembedding the current hypothesis. The embedding vector enters only through AdaLN modulation, making the approach encoder agnostic: the same architecture applies to any embedding model without alignment training. We validate on three encoders with different architectures and dimensionalities, with no access to the target encoder at inference time.
@@ 3.3 Model Architecture
where $\mathbf{W}_{1}\in\mathbb{R}^{D_{h}\times d}$, $\mathbf{W}_{2}\in\mathbb{R}^{D_{h}\times D_{h}}$, and $\mathbf{b}_{1},\mathbf{b}_{2}\in\mathbb{R}^{D_{h}}$ are learned parameters. We primarily use a 22-layer transformer initialized from multilingual BERT with vocabulary size 256K. With $D_{h}=768$ and FFN dimension 3072, the model has 388M total parameters. When freezing the pretrained backbone, 191M parameters remain trainable (49.3%), consisting of the embedding projection MLP, AdaLN conditioning MLPs, and output layer. We also experiment with 8-layer (268M parameters) and 2-layer (217M parameters with 20M trainable when frozen) configurations to assess depth-accuracy tradeoffs. Input and output embeddings are weight-tied to reduce parameters given the large vocabulary size.
@@
where $\text{MLP}_{t}^{(\ell)}$ and $\text{MLP}_{c}^{(\ell)}$ are single-layer MLPs that output vectors of dimension $D_{h}$, $\gamma^{(\ell)}$ is the scale parameter, and $\beta^{(\ell)}$ is the shift parameter. The layer normalization at layer $\ell$ is then modulated:
@@ 3.4 Decoding Strategies
We consider five strategies for generating tokens from the trained model.
@@
Confidence-based decoding adapts the MaskGIT approach to text generation. At each iteration step $s$, we predict all token positions and unmask the top-$k_{s}$ most confident tokens:
$c_{i}$ $=\max_{v\in\mathcal{V}}p_{\theta}(v\mid x_{t},t,\mathbf{e})\quad\forall\,i$ (14)
$x_{t^{\prime},i}$ $=\begin{cases}\arg\max_{v\in\mathcal{V}}p_{\theta}(v\mid x_{t},t,\mathbf{e})&\text{if }c_{i}\text{ is in top-}k_{s}\text{ confidence scores}\\ [\text{MASK}]&\text{otherwise}\end{cases}$ (15)
where $c_{i}$ is the confidence at position $i$, and $k_{s}$ decreases from $n$ to 0 over a fixed number of iterations. This allows the model to first commit to high-confidence predictions and progressively fill uncertain positions using the context from already decoded tokens.
@@ 4 Experimental Results
We train on 2M samples from C4 (Raffel et al., 2020), filtered to 32 tokens. Training uses batch size 380-400 for up to 200K steps with AdamW optimizer at learning rate $10^{-4}$, warmup of 2000 steps, and EMA decay 0.9999. We employ a log-linear noise schedule with $\lambda=5.0$ following Sahoo et al. (2024). Timesteps are sampled uniformly from $[0,1]$. Embeddings are computed using the target encoder and cached. We evaluate on three embedding models with different architectures and dimensionalities: jina-embeddings-v3 (Sturua et al., 2025) with 570M parameters and 1024-dimensional embeddings, Qwen3-Embedding-0.6B with 600M parameters and 1024-dimensional embeddings, and EmbeddingGemma-300m with 300M parameters and 768-dimensional embeddings. We train separate models for each encoder using multilingual data from mC4 to assess generalization across embedding spaces.
The 22-layer configuration with 256K vocabulary has 388M total parameters. When freezing the pretrained mmBERT backbone, 191M parameters (49.3%) remain trainable, consisting of the embedding projection MLP, AdaLN conditioning MLPs, and output layer. The 8-layer configuration has 268M total parameters, all trainable. The 2-layer configuration has 217M total parameters with 20M trainable when frozen.
Table 1 shows results across all three embedding encoders using sequential greedy decoding. These numbers represent training accuracy on the cached embedding dataset and may not reflect generalization to held-out test data. The reported accuracies vary by encoder: Qwen3-Embedding at 81.3%, EmbeddingGemma at 78.8%, and jina-v3 at 76.0%. All models are trained on multilingual data from mC4. Current ablation experiments with the 22-layer mmBERT backbone show training accuracies of 68.2% for the no-freeze configuration and 63.8% for the freeze configuration, indicating substantial variance across experimental setups and checkpoints.
Table 1: Performance across embedding encoders using sequential greedy decoding. All trained on 2M multilingual samples from mC4. Accuracies are training metrics on cached embeddings and should not be interpreted as test-set performance. Best checkpoint selected by validation loss.
@@
Table 2 compares decoding strategies across all three encoders on 10 languages. Cosine similarity is averaged over the same sentence translated into English, Chinese, German, Japanese, French, Spanish, Korean, Russian, Arabic, and Portuguese. Sequential greedy decoding shows the highest similarity for jina-v3 and EmbeddingGemma, while two-stage decoding performs best for Qwen3-Embedding.
@@
Euler with remasking at 0.05 improves over vanilla Euler by 2.6 percentage points in token accuracy. Two-stage decoding achieves the highest exact match rate at 13.1%. Baselines confirm that embedding conditioning is necessary: random tokens achieve 0.02% accuracy, while an unconditional language model achieves 2.1% accuracy despite high fluency with BLEU score 89.3, indicating that fluency alone does not ensure faithful reconstruction.
@@
Figure 2: Training dynamics across three embedding encoders on 2M multilingual samples. Qwen3-Embedding reaches 81.3% training token accuracy at 72.5K steps with validation loss 1.32. All models show plateauing behavior beyond 50K steps, indicating that further improvements may require architectural changes rather than extended training.
5 Conclusion and Future Work
We presented embedding inversion via conditional masked diffusion, demonstrating that parallel denoising can recover tokens from embedding vectors without access to the target encoder. The method operates in a strictly black-box setting: the embedding vector is the only input at inference time, with no correction loop through the encoder. This threat model is more restrictive than Vec2Text, which requires encoder access for iterative refinement, or ALGEN, which needs alignment training on encoder outputs.
This architectural difference also explains the performance gap. Autoregressive methods with iterative correction benefit from a closed-loop feedback signal. Each refinement step re-embeds the current hypothesis and computes an explicit error signal against the target. Diffusion-based inversion lacks this mechanism. The embedding conditions the generation through AdaLN modulation but never verifies whether the output actually maps back to the target embedding. Errors compound without correction, which our ablation results confirm: dynamic masking schedules that provide implicit curriculum learning substantially outperform fixed schedules, partially compensating for the absence of explicit feedback.
Two directions appear promising for closing this gap. Classifier-free guidance could inject stronger embedding signal during sampling without requiring encoder access at inference time. Alternatively, combining diffusion-based initialization with a lightweight correction step through a distilled or approximate encoder could recover much of the closed-loop advantage while maintaining broader applicability across encoder architectures. Scaling to longer sequences via hierarchical diffusion over sentence-level chunks is a natural next step.
Appendix A Mask Ratio Ablation with mmBERT Decoder
We conduct a separate ablation study using a different decoder architecture to investigate the effect of fixed mask ratios on training dynamics. Unlike the main experiments, which use an 8-layer transformer trained from scratch with the GPT-2 tokenizer, these experiments use a 22-layer transformer initialized from mmBERT pretrained weights with a 256K multilingual vocabulary. Five models were trained with fixed mask percentages of 10%, 20%, 40%, 80%, and 100%, sharing identical architecture (388M parameters), training configuration (batch size 380, learning rate $10^{-4}$, warmup 2000 steps), and pretrained initialization.
Table 4: Mask ratio ablation with mmBERT decoder. All models early-stopped before 200K steps.
Mask % Best Step Best Val Loss Final Train Acc Tokens (M)
10% 43,500 5.9943 30.2% 3,413
20% 42,500 5.9197 21.0% 3,284
40% 38,500 5.8730 14.0% 3,174
80% 35,500 5.7933 10.8% 2,955
100% 30,500 5.8410 10.1% 2,590
Higher mask ratios achieve lower validation loss, with 80% masking reaching 5.79 (Table 4). Training accuracy inversely correlates with mask ratio as expected. Higher mask ratios also converge faster, with 100% masking reaching its best validation loss at step 30,500 versus 43,500 for 10% masking. However, all fixed-mask models underperform configurations using the log-linear schedule from the main experiments, suggesting that dynamic masking provides curriculum learning benefits.
(a) Training accuracy curves
(b) Validation loss curves
Figure 3: Training dynamics across fixed mask ratios with mmBERT decoder. All models plateau early, and higher training accuracy from lower mask ratios does not translate into better validation performance.